# Second-Order Fine-Tuning without Pain for LLMs: A Hessian Informed Zeroth-Order Optimizer

### v3

@@ 4.3 Training with Non-Differentiable Objectives
Our proposed HiZOO employs gradient estimation to update parameters, allowing for the use of non-differentiable objectives for training. Following the setting of MeZO [32], we conduct extensive experiments using F1 as optimization objective. The results presented in Table 4 indicate that our method outperforms MeZO by 6.36% on F1 on average.
@@ 4.4 Memory Usage and Time Efficiency Analysis
[…]
Time Efficiency    In Appendix F, we analyse the wall-clock time efficiencies and show that HiZOO spends 1.5×1.5$\times$1.5 × time per step compared with MeZO, mainly from the extra forward pass. However, HiZOO reduces total number of forward passes required for convergence. For example, HiZOO achieves a 8×8$\times$8 × and 4×4$\times$4 × speedup on SST2 and MNLI tasks.
@@ 4.5 Comparison with other ZO variants
We also compare our HiZOO with a broader array of ZO optimization techniques [59]. As shown in Table 5, our HiZOO outperforms all other ZO methods. Compared with ZO-Adam who leverages second-order moment to guide gradient descent, our HiZOO-L achieves a notable 4.3% absolute improvement, while using 50% of the GPU memory.
@@ 4.6 Hyperparameter Analysis
Influence of Smooth Scale $\alpha_{t}$ in EMA    To assess the robustness of the optimizer, a grid search is conducted to evaluate the sensitivity of the hyper-parameter $\alpha_{t}$ on RoBERTa-large (350M). Figure 7 illustrates that the smoothing scale $\alpha_{t}$ is instrumental in controlling the convergence rate during training. As $\alpha_{t}$ is incrementally increased from zero, the training loss decreases faster. However, too large $\alpha_{t}$ values may impede convergence or even cause training to fail due to gradient explosion.
@@
Figure 6: Experiments on influence of EMA $\alpha_{t}$ for hessian in Eq. (6). We use HiZOO (prefix) to fine-tune Roberta-large on SNLI. More results can be found in Appendix H.1.
Figure 7: Loss curves on Function (a) using the variant HiZOO-multi with different estimation times $n$ per step. Trajectory visualization can be found in Appendix D.1.
@@ 5 Conclusion
In this work, we introduce HiZOO, which is the first ZOO that incorporates diagonal Hessian for fine-tuning LLMs. By introducing one more forward pass, HiZOO can handle heterogeneous curvatures across different parameter dimensions. We provide theoretical analysis and visualize the optimization trajectories to explore how it works. Further experiments on LLMs show that HiZOO converges in much fewer steps than MeZO and achieves better performance, across various models (RoBERTa-large, OPT, Llama3, Phi-2) and methods (FT, LoRA and prefix). In addition, HiZOO increases extra memory cost because of the storage of Hessian. We explore a memory efficient implementation (HiZOO-L) and there are many other methods worth trying, such as quantization [10], FlashAttention [8].
@@ A.1 First-order Optimizer Used in LLMs
Optimization methods have consistently been a popular research domain, encompassing techniques such as Gradient Descent (GD), Momentum, Adagrad [12], ADADELTA [54], and Newton’s method, which have been instrumental in advancing fields like computer vision. However, the emergence of large-scale models, characterized by their massive parameter counts and intricate architectures, has challenged the efficacy of conventional optimization methods for training tasks. Amidst this landscape, Adam [22] has emerged as the preferred choice for its ability to rapidly converge, making it particularly suitable for the training and fine-tuning large models. Then AdamW [29] was proposed to add a weight decay coefficient to alleviate over-fitting. Notwithstanding these advancements, a limitation persists with these optimizers: they have an implicit batch size ceiling. Exceeding this threshold can provoke extreme gradient updates, thus impeding the convergence rate of the models. This bottleneck is particularly problematic in the context of large-model training, which typically necessitates substantial batch sizes. To circumvent this constraint, LAMB [53] was devised to apply principled layer-wise adaptation strategy to accelerate the training of large models employing large batches.
@@ A.2 Hessian Based First-Order Optimizer
Compared with first-order optimizers, second-order optimizer considers second-order information in the process of gradient calculation. As a result, it has more abundant information to guide gradient descent and is considered to be more promising. Previous studies utilized curvature information to pre-condition the gradient [4, 34, 7]. Subsequently, Magoulas et al. [31] applied diagonal Hessian as the pre-conditioner, which greatly promotes the landing of second-order optimizer in the field of deep learning. Martens [33] approximated the Hessian with conjugate gradient. Schaul et al. [40] utilized diagonal Hessian to automatically adjust the learning rate of SGD during training. Another work [36] extended natural gradient descent to incorporate second order information alongside the manifold information and used a truncated Newton approach for inverting the metric matrix instead of using a diagonal approximation of it. EVA  [56] proposed to use the Kronecker factorization of two small vectors to approximated the Hessian, which significantly reduces memory consumption. AdaHessian [50] incorporates an approximate Hessian diagonal, with spatial averaging and momentum to precondition the gradient vector.
Although great progress has been made in the research of second-order optimizer, it has not been widely used because of the extra computation and memory cost when gradient updating, and this situation is extremely serious in the training of large language models. Based on the above dilemma, recent works [2, 14] proposed to offload Hessian computation to CPUs and utilized ResNets and very large batch size to approximate the Fisher information matrix. Sophia [26] was the first to apply second-order optimizer and achieve a speed-up on large language models in total compute successfully.
@@ A.3 Zeroth-Order Optimizer
Zeroth-order optimization, is also known as derivative-free or black-box optimization. There have been many one-point gradient estimators in past works [13, 43, 46, 42, 20, 1, 38, 47]. However, cursory experiments with one such promising estimator [58] reveal that SPSA outperforms other methods.
In previous works, it appears in a wide range of applications where either the objective function is implicit or its gradient is impossible or too expensive to compute. For example, methods [44, 18] consider derivative-free distributed algorithms for non-convex multi-agent optimization. ZO-BCD[5], ZOO[6], ZO-signSGD [27] and ZO-HessAware [52] utilize zeroth-order stochastic optimization to generate black-box adversarial example in deep learning.
Beyond that, MeZO [32] firstly adapted the classical ZO-SGD method to fine-tune LLMs, while achieving comparable performance with extremely great memory reduction and GPU-hour reduction. Subsequently, ZO-AdaMU [21] improved ZO-SGD and adapted the simulated perturbation with momentum in its stochastic approximation method. Both of these two optimizers provide researchers with a new and promising technique for fine-tuning large models.
@@ Appendix B Detailed Convergence Analysis
The objective function $\mathcal{L}(\theta;\mathcal{B})$ is $L$-smooth, which means that for any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, it holds that:
$\mathcal{L}(\theta_{2};\mathcal{B})\leq\mathcal{L}(\theta_{1};\mathcal{B})+\langle\nabla\mathcal{L}(\theta_{1};\mathcal{B}),\theta_{2}-\theta_{1}\rangle+\frac{L}{2}\|\theta_{2}-\theta_{1}\|^{2}.$ (10)
@@
The stochastic gradient $\nabla\mathcal{L}(\theta;\mathcal{B})$ has $\sigma^{2}$ variance, which means:
$\mathbb{E}\left[\|\nabla\mathcal{L}(\theta;\mathcal{B})-\nabla\mathcal{L}(\theta)\|^{2}\right]\leq\sigma^{2}.$ (11)
@@
$\mathbb{E}\left[\mathcal{L}(\theta_{t+1};\mathcal{B}_{t+1})\mid\theta_{t}\right]$
$\leq$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\eta_{t}\mathbb{E}\left[\langle\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t}),g_{\mu}(\theta_{t})\rangle\right]+\frac{L\eta_{t}^{2}}{2}\mathbb{E}\left[\|g_{\mu}(\theta_{t})\|^{2}\right]$
$\leq$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\eta_{t}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}+\eta_{t}\mathcal{O}\left(\mu\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|\right)$
$+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}$
@@
$\leq$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\frac{\eta_{t}}{2}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}$
@@
$=$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\frac{\eta_{t}}{2}\left(1-4\eta_{t}L(\mathrm{tr}(\Sigma_{t})+\beta_{u})\right)\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}$
@@
$\leq$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\frac{\eta_{t}}{4}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(\mu^{2}),$
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\frac{\eta_{t}}{4}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}\right]\leq\sum_{t=1}^{T}\left(\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\mathcal{L}(\theta_{t+1};\mathcal{B}_{t+1})\right)$
@@
$=$ $\mathcal{L}(\theta_{1};\mathcal{B}_{1})-\mathcal{L}(\theta_{T+1};\mathcal{B}_{T+1})+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(T\mu^{2})$
$\leq$ $\mathcal{L}(\theta_{1};\mathcal{B}_{1})-\mathcal{L}(\theta_{*};\mathcal{B}_{*})+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(T\mu^{2}).$
@@
$\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{\mbox{out}};\mathcal{B})\|^{2}\right]=\frac{1}{T}\sum_{t=1}^{T}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B})\|^{2}\leq\frac{1}{T\beta_{\ell}}\sum_{t=1}^{T}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B})\|_{\Sigma_{t}}^{2}$
$\leq$ $\frac{4(\mathcal{L}(\theta_{1};\mathcal{B})-\mathcal{L}(\theta_{*};\mathcal{B}))}{T\beta_{\ell}\eta}+\frac{8\eta L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)}{T\beta_{\ell}}\sigma^{2}+\mathcal{O}(\mu^{2})$
$=$ $\frac{32L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)(\mathcal{L}(\theta_{1};\mathcal{B})-\mathcal{L}(\theta_{*};\mathcal{B}))}{\sqrt{T}\beta_{\ell}}+\frac{\sigma^{2}}{T^{3/2}\beta_{\ell}}+\mathcal{O}\left(\mu^{2}\right),$
@@
Eq. (8) shows that once we choose the step size $\eta$ properly, $\mathcal{L}(\theta_{t+1};\mathcal{B}_{t+1})$ will be less than $\mathcal{L}(\theta_{t};\mathcal{B}_{t})$ in expectation up to some noises of order $\mu^{2}$. Specifically, if set $\eta=\frac{1}{8\sqrt{T}L(\max_{t}\mathrm{tr}(\Sigma_{t})+\beta_{u})}$, Eq. (9) implies that we can find an solution $\theta_{\mbox{out}}$ such that $\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{\mbox{out}};\mathcal{B})\|^{2}\right]\leq\epsilon^{2}$ in $\mathcal{O}(\epsilon^{-4})$ iterations. This rate matches the one of [15].
@@
We assume that Assumption B.2 and Assumption B.3 hold. Then, $g_{\mu}(\theta_{t})$ defined in Eq. (7) has the following properties:
$\mathbb{E}\left[g_{\mu}(\theta_{t})\right]=\Sigma_{t}\nabla\mathcal{L}(\theta_{t})+\mathcal{O}(\mu)$
$\mathbb{E}\left[\|g_{\mu}(\theta_{t})\|^{2}\right]\leq 4\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}$
@@
$=$ $\sum_{i=1}^{b}\frac{\mathcal{L}(\theta_{t}+\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B}_{t})-\mathcal{L}(\theta_{t}-\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B}_{t})}{2b\mu}\Sigma_{t}^{1/2}u_{i}$
$=$ $\sum_{i=1}^{b}\frac{2\mu\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}^{1/2}u_{i}+\mathcal{O}(\mu^{2})}{2b\mu}\Sigma_{t}^{1/2}u_{i}$
$=$ $\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu).$
@@
$\mathbb{E}\left[g_{\mu}(\theta_{t})\right]=\Sigma_{t}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu).$ (12)
@@
$=$ $\mathbb{E}\bigg{[}\|\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu)\|^{2}\bigg{]}$
$\leq$ $2\mathbb{E}\left[\|\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|^{2}\right]+\mathcal{O}(\mu^{2})$
$\leq$ $\frac{2}{b}\sum_{i=1}^{b}\mathbb{E}\left[\|\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|^{2}\right]+\mathcal{O}(\mu^{2})$
$=$ $2\mathrm{tr}(\Sigma_{t})\cdot\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})$
$+2\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}^{2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu^{2})$
$\leq$ $2\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu^{2}),$

### v4

@@ 4.3 Training with Non-Differentiable Objectives
Our proposed HiZOO employs gradient estimation to update parameters, allowing for the use of non-differentiable objectives for training. Following the setting of MeZO (Malladi et al., 2023), we conduct extensive experiments using F1 as optimization objective. The results presented in Table 4 indicate that our method outperforms MeZO by 6.54% absolute on F1 on average.
@@ 4.4 Memory Usage and Time Efficiency Analysis
[…]
Time Efficiency    We analyse the wall-clock time efficiencies and find that HiZOO and HiZOO-L spend 1.5×1.5$\times$1.5 × time per step compared with MeZO, mainly from the extra forward pass, details in Appendix G. However, HiZOO reduces total number of forward passes required for convergence. For example, HiZOO achieves a 8$\times$ and 4$\times$ speedup on SST2 and MNLI tasks.
@@ 4.5 Comparison with other ZO variants
We also compare our HiZOO with a broader array of ZO optimization techniques Zhang et al. (2024). As shown in Table 5, our HiZOO outperforms all other ZO methods. Compared with ZO-Adam who leverages second-order moment to guide gradient descent, our HiZOO-L achieves a notable 4.3% absolute improvement, while using 50% of the GPU memory.
@@ 4.6 Hyperparameter Analysis
Figure 6: Influence of EMA $\alpha_{t}$ for hessian in Eq. equation 5. We use HiZOO (prefix) to fine-tune Roberta-large on SNLI. More results can be found in Appendix H.1.
Figure 7: Loss curves on Function (a) using the variant HiZOO-multi with different estimation times $n$ per step. Trajectory visualization can be found in Appendix H.1.
Influence of Smooth Scale $\alpha_{t}$ in EMA    To assess the robustness of the optimizer, a grid search is conducted to evaluate the sensitivity of the hyper-parameter $\alpha_{t}$ on RoBERTa-large (350M). Figure 7 illustrates that as $\alpha_{t}$ is incrementally increased from zero, the training loss decreases faster. However, too large $\alpha_{t}$ values may impede convergence or even cause training to fail due to gradient explosion.
@@ 5 Conclusion
In this work, we introduce HiZOO, which is the first ZOO that incorporates diagonal Hessian for fine-tuning LLMs. By introducing one more forward pass, HiZOO can handle heterogeneous curvatures across different parameter dimensions. We provide theoretical analysis and visualize the optimization trajectories to explore how it works. Further experiments show that HiZOO converges in much fewer steps than MeZO and achieves better performance across various LLMs. We also explore a memory efficient implementation (HiZOO-L) to reduce the Hessian consumption.
@@ A.1 First-order Optimizer Used in LLMs
Optimization methods have consistently been a popular research domain, encompassing techniques such as Gradient Descent (GD), Momentum, Adagrad (Duchi et al., 2011), ADADELTA (Zeiler, 2012), and Newton’s method, which have been instrumental in advancing fields like computer vision. However, the emergence of large-scale models, characterized by their massive parameter counts and intricate architectures, has challenged the efficacy of conventional optimization methods for training tasks. Amidst this landscape, Adam (Kingma & Ba, 2015) has emerged as the preferred choice for its ability to rapidly converge, making it particularly suitable for the training and fine-tuning large models. Then AdamW (Loshchilov & Hutter, 2019) was proposed to add a weight decay coefficient to alleviate over-fitting. Notwithstanding these advancements, a limitation persists with these optimizers: they have an implicit batch size ceiling. Exceeding this threshold can provoke extreme gradient updates, thus impeding the convergence rate of the models. This bottleneck is particularly problematic in the context of large-model training, which typically necessitates substantial batch sizes. To circumvent this constraint, LAMB (You et al., 2020) was devised to apply principled layer-wise adaptation strategy to accelerate the training of large models employing large batches.
@@ A.2 Hessian Based First-Order Optimizer
Compared with first-order optimizers, second-order optimizer considers second-order information in the process of gradient calculation. As a result, it has more abundant information to guide gradient descent and is considered to be more promising. Previous studies utilized curvature information to pre-condition the gradient (BROYDEN, 1970; Nesterov & Polyak, 2006; Conn et al., 2000). Subsequently, Magoulas et al. (Magoulas et al., 1999) applied diagonal Hessian as the pre-conditioner, which greatly promotes the landing of second-order optimizer in the field of deep learning. Martens (Martens, 2010) approximated the Hessian with conjugate gradient. Schaul et al. (Schaul et al., 2013) utilized diagonal Hessian to automatically adjust the learning rate of SGD during training. Another work (Pascanu & Bengio, 2014) extended natural gradient descent to incorporate second order information alongside the manifold information and used a truncated Newton approach for inverting the metric matrix instead of using a diagonal approximation of it. EVA  (Zhang et al., 2023) proposed to use the Kronecker factorization of two small vectors to approximated the Hessian, which significantly reduces memory consumption. AdaHessian (Yao et al., 2021) incorporates an approximate Hessian diagonal, with spatial averaging and momentum to precondition the gradient vector.
Although great progress has been made in the research of second-order optimizer, it has not been widely used because of the extra computation and memory cost when gradient updating, and this situation is extremely serious in the training of large language models. Based on the above dilemma, recent works (Anil et al., 2021; George et al., 2018) proposed to offload Hessian computation to CPUs and utilized ResNets and very large batch size to approximate the Fisher information matrix. Sophia (Liu & Li, 2023) was the first to apply second-order optimizer and achieve a speed-up on large language models in total compute successfully.
@@ A.3 Zeroth-Order Optimizer
Zeroth-order optimization, is also known as derivative-free or black-box optimization. There have been many one-point gradient estimators in past works (FairScale authors, 2021; Spall, 1997; Vakhitov et al., 2009; Spall, 1992; Jamieson et al., 2012; Agarwal et al., 2009; Raginsky & Rakhlin, 2011; Wang et al., 2020). However, cursory experiments with one such promising estimator (Zhang et al., 2022b) reveal that SPSA outperforms other methods.
In previous works, it appears in a wide range of applications where either the objective function is implicit or its gradient is impossible or too expensive to compute. For example, methods (Tang et al., 2021; Hajinezhad & Zavlanos, 2018) consider derivative-free distributed algorithms for non-convex multi-agent optimization. ZO-BCD(Cai et al., 2021), ZOO(Chen et al., 2017), ZO-signSGD (Liu et al., 2019a) and ZO-HessAware (Ye et al., 2019) utilize zeroth-order stochastic optimization to generate black-box adversarial example in deep learning.
Beyond that, MeZO (Malladi et al., 2023) firstly adapted the classical ZO-SGD method to fine-tune LLMs, while achieving comparable performance with extremely great memory reduction and GPU-hour reduction. Recently there have been many subsequent excellent works (Jiang et al., 2023; Zhao et al., 2024a; Liu et al., 2024; Guo et al., 2024; Tang, 2024; Chen et al., 2024; Wang et al., ; Chen et al., 2025a; Tan et al., 2025; Chen et al., 2025b; Sun et al., 2025). All of these optimizers provide researchers with a new and promising technique for fine-tuning large models.
@@ Appendix B Detailed Convergence Analysis
The objective function ${\bm{L}}(\theta)$ is $L$-smooth, which means that for any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, it holds that:
${\bm{L}}(\theta_{2})\leq{\bm{L}}(\theta_{1})+\langle\nabla{\bm{L}}(\theta_{1}),\theta_{2}-\theta_{1}\rangle+\frac{L}{2}\|\theta_{2}-\theta_{1}\|^{2}.$ (9)
@@
The stochastic gradient $\nabla{\bm{L}}(\theta)$ has $\sigma^{2}$ variance, which means:
$\mathbb{E}\left[\|\nabla{\bm{L}}(\theta)-\nabla{\bm{L}}(\theta)\|^{2}\right]\leq\sigma^{2}.$ (10)
@@
$\mathbb{E}\left[{\bm{L}}(\theta_{t+1})\mid\theta_{t}\right]$
$\leq$ ${\bm{L}}(\theta_{t})-\eta_{t}\mathbb{E}\left[\langle\nabla{\bm{L}}(\theta_{t}),g_{\mu}(\theta_{t})\rangle\right]+\frac{L\eta_{t}^{2}}{2}\mathbb{E}\left[\|g_{\mu}(\theta_{t})\|^{2}\right]$
$\leq$ ${\bm{L}}(\theta_{t})-\eta_{t}\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}+\eta_{t}\mathcal{O}\left(\mu\|\nabla{\bm{L}}(\theta_{t})\|\right)$
$+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}$
@@
$\leq$ ${\bm{L}}(\theta_{t})-\frac{\eta_{t}}{2}\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}$
@@
$=$ ${\bm{L}}(\theta_{t})-\frac{\eta_{t}}{2}\left(1-4\eta_{t}L(\mathrm{tr}(\Sigma_{t})+\beta_{u})\right)\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}$
@@
$\leq$ ${\bm{L}}(\theta_{t})-\frac{\eta_{t}}{4}\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(\mu^{2}),$
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\frac{\eta_{t}}{4}\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]\leq\sum_{t=1}^{T}\left({\bm{L}}(\theta_{t})-{\bm{L}}(\theta_{t+1})\right)$
@@
$=$ ${\bm{L}}(\theta_{1})-{\bm{L}}(\theta_{T+1})+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(T\mu^{2})$
$\leq$ ${\bm{L}}(\theta_{1})-{\bm{L}}(\theta_{*})+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(T\mu^{2}).$
@@
$\mathbb{E}\left[\|\nabla{\bm{L}}(\theta_{\mbox{out}})\|^{2}\right]=\frac{1}{T}\sum_{t=1}^{T}\|\nabla{\bm{L}}(\theta_{t})\|^{2}\leq\frac{1}{T\beta_{\ell}}\sum_{t=1}^{T}\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}$
$\leq$ $\frac{4({\bm{L}}(\theta_{1})-{\bm{L}}(\theta_{*}))}{T\beta_{\ell}\eta}+\frac{8\eta L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)}{T\beta_{\ell}}\sigma^{2}+\mathcal{O}(\mu^{2})$
$=$ $\frac{32L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)({\bm{L}}(\theta_{1})-{\bm{L}}(\theta_{*}))}{\sqrt{T}\beta_{\ell}}+\frac{\sigma^{2}}{T^{3/2}\beta_{\ell}}+\mathcal{O}\left(\mu^{2}\right),$
@@
Eq. equation 7 shows that once we choose the step size $\eta$ properly, ${\bm{L}}(\theta_{t+1})$ will be less than ${\bm{L}}(\theta_{t})$ in expectation up to some noises of order $\mu^{2}$. Specifically, if set $\eta=\frac{1}{8\sqrt{T}L(\max_{t}\mathrm{tr}(\Sigma_{t})+\beta_{u})}$, Eq. equation 8 implies that we can find an solution $\theta_{\mbox{out}}$ such that $\mathbb{E}\left[\|\nabla{\bm{L}}(\theta_{\mbox{out}})\|^{2}\right]\leq\epsilon^{2}$ in $\mathcal{O}(\epsilon^{-4})$ iterations. This rate matches the one of (Ghadimi & Lan, 2013).
@@
We assume that Assumption B.2 and Assumption B.3 hold. Then, $g_{\mu}(\theta_{t})$ defined in Eq. equation 6 has the following properties:
$\mathbb{E}\left[g_{\mu}(\theta_{t})\right]=\Sigma_{t}\nabla{\bm{L}}(\theta_{t})+\mathcal{O}(\mu)$
$\mathbb{E}\left[\|g_{\mu}(\theta_{t})\|^{2}\right]\leq 4\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}$
@@
$=$ $\sum_{i=1}^{b}\frac{{\bm{L}}(\theta_{t}+\mu\Sigma_{t}^{1/2}u_{i})-{\bm{L}}(\theta_{t}-\mu\Sigma_{t}^{1/2}u_{i})}{2b\mu}\Sigma_{t}^{1/2}u_{i}$
$=$ $\sum_{i=1}^{b}\frac{2\mu\nabla^{\top}{\bm{L}}(\theta_{t})\Sigma_{t}^{1/2}u_{i}+\mathcal{O}(\mu^{2})}{2b\mu}\Sigma_{t}^{1/2}u_{i}$
$=$ $\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla{\bm{L}}(\theta_{t})+\mathcal{O}(\mu).$
@@
$\mathbb{E}\left[g_{\mu}(\theta_{t})\right]=\Sigma_{t}\nabla{\bm{L}}(\theta_{t})+\mathcal{O}(\mu).$ (11)
@@
$=$ $\mathbb{E}\bigg{[}\|\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla{\bm{L}}(\theta_{t})+\mathcal{O}(\mu)\|^{2}\bigg{]}$
$\leq$ $2\mathbb{E}\left[\|\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla{\bm{L}}(\theta_{t})\|^{2}\right]+\mathcal{O}(\mu^{2})$
$\leq$ $\frac{2}{b}\sum_{i=1}^{b}\mathbb{E}\left[\|\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla{\bm{L}}(\theta_{t})\|^{2}\right]+\mathcal{O}(\mu^{2})$
$=$ $2\mathrm{tr}(\Sigma_{t})\cdot\nabla^{\top}{\bm{L}}(\theta_{t})\Sigma_{t}\nabla{\bm{L}}(\theta_{t})$
$+2\nabla^{\top}{\bm{L}}(\theta_{t})\Sigma_{t}^{2}\nabla{\bm{L}}(\theta_{t})+\mathcal{O}(\mu^{2})$
$\leq$ $2\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\nabla^{\top}{\bm{L}}(\theta_{t})\Sigma_{t}\nabla{\bm{L}}(\theta_{t})+\mathcal{O}(\mu^{2}),$